A sequence of heads(H) and tails (T) in tossing of a coin 16 times are given below HTTTHTHTHHTHTTHH count the number of runs and also test whether the heads and the tails occur in a random order alpha =0.05
1
Expert's answer
2021-08-05T13:48:48-0400

For the given sequence, the sample size "n=16," number of heads "n_1=8," number of tails "n_2=8."


"(H)(TTT)(H)(T)(H)(T)(HH)(T)(H)(TT)(HH)"

number of runs of "H, R_1=6"

number of runs of "T, R_2=5"

So that number of runs is "R=R_1+R_2=6+5=11."

"H_0:" the sequence was produced in a random manner.

"H_1:" the sequence was not produced in a random manner.



"\mu=\dfrac{2n_1n_2}{n_1+n_2}+1=\dfrac{2(8)(8)}{8+8}+1=9"

"s^2=\dfrac{2n_1n_2(2n_1n_2-n_1-n_2)}{(n_1+n_2)^2(n_1+n_2-1)}"

"=\dfrac{2(8)(8)(2(8)(8)-8-8)}{(8+8)^2(8+8-1)}=\dfrac{56}{15}"

"z=\dfrac{R-\mu}{s}=\dfrac{11-9}{\sqrt{\dfrac{56}{15}}}\approx1.0351"

For "\alpha=0.05," "z_c=1.96."

Since "z=1.0351<1.96=z_c," we accept "H_0." It means that the heads and tails occur in random order or it can be said that the coin is unbiased at the "\alpha=0.0" significance level.